OpenStudy (anonymous):

I I am still trying to work the problem m^3-6m/m^2-64=?

OpenStudy (anonymous):

\[\frac{m^3-6m}{m^2-64}=?\]is this it?

OpenStudy (anonymous):

What I would do is factor the bottom first and get \[m^2-64=(m-8)(m+8)\]and try to factor the top out to get one those but can't seem to figure it out. Anyone care to jump in?

OpenStudy (anonymous):

And you are just trying to simplify this right?